# Bluesky and the AT Protocol: Usable Decentralized Social Media

### v1

@@ Abstract.
Bluesky is a new social network built upon the AT Protocol, a decentralized foundation for public social media. It was launched in private beta in February 2023, and has grown to over 3 million registered users in the following year. In this paper we introduce the architecture of Bluesky and the AT Protocol, which is inspired by the web itself, but modernized to include streams of real-time updates and cryptographic authentication. We explain how the technical design of Bluesky is informed by our goals: to enable decentralization by having multiple interoperable providers for every part of the system; to make it easy for users to switch providers; to give users agency over the content they see; and to provide a simple user experience that does not burden users with complexity arising from the system’s decentralized nature. The system’s openness allows anybody to contribute to content moderation and community management, and we invite the research community to use Bluesky as a dataset and testing ground for new approaches in social media moderation.
copyright: none
@@ 1. Introduction
Over the last two decades, social media services have evolved from a fun curiosity into a cornerstone of civic life (Barabas et al., 2017). This development has been accompanied by increasing unease that mainstream “digital town squares”, such as Twitter/X or Facebook, are under the control of a single corporation, and may change their policies on the whim of their leaders (Yeung, 2023). Their operations are opaque (e.g. regarding which content is recommended to users), and their users lack agency over their user experience. As a result, there has been increasing interest in decentralized social networks, of which the fediverse around the ActivityPub protocol (Lemmer-Webber et al., 2018) and the Mastodon software (Mastodon gGmbH, [n. d.]) is perhaps the best known (we review a selection of decentralized social networks in Section 4).
@@
In this paper we introduce the AT Protocol (atproto), a decentralized foundation for social networking, and Bluesky, a Twitter-style social app built upon it. A core design goal of atproto and Bluesky is to enable a user experience of the same or better quality as centralized services, while being open and decentralized on a technical level. We introduce the user-facing features of Bluesky in Section 2, and in Section 3 we explain the underlying systems architecture. The AT Protocol is designed such that for every part of the system there are multiple competing operators providing interoperable services, making it easy to switch from one provider to another.
Decentralization alone is not able to solve some of the thorniest problems of social media, such as misinformation, harassment, and hate speech (Roth and Lai, 2023). However, by opening up the internals of a service to contributors who are not employees of a particular company, decentralization can enable a marketplace of approaches to these problems (Masnick, 2019). For example, Bluesky allows anybody to run moderation services that make subjective decisions of selecting desirable content or flagging undesirable content, and users can choose which moderation services they want to subscribe to. Moderation services are decoupled from hosting providers, making it easy for users to switch moderation services until they find ones that match their preferences. Our hope is that this architectural openness enables communities to develop their own approaches to managing problematic content, independently of what any particular service operator implements (Masnick, 2019).
For example, researchers wanting to identify disinformation campaigns can easily get access to all content being posted, the social graph, and user profiles on Bluesky. If they are able construct an algorithm to label suspected disinformation, they can publish their labels in real time, and users who wish to see those labels can enable them in their client software. One goal of this paper is to bring Bluesky and the AT Protocol to the attention of researchers working on such algorithms, and to invite them to use the rapidly growing dataset of Bluesky content as a basis for their work.
Refer to caption A Twitter-like feed of short posts. At the top is a feed selector, in which the default “Following” feed is active.
Figure 1. Screenshot of the Bluesky home screen.
@@ 2. The Bluesky Social App
Bluesky presents itself to users as a straightforward microblogging application in the style of Twitter/X (see Figure 1). The “official” client app is available on iOS, Android, and the web; several independently developed client apps are also available, such as Graysky (Newman, [n. d.]) and deck.blue (Filho, [n. d.]). Users can make public posts containing up to 300 characters of text, and up to four images, and they can interact with posts by replying, reposting, or liking. A user can also follow other users, and the default feed shows posts by accounts that the user is following in reverse chronological order. There are also alternative feeds that show content on various topics, without the user needing to follow the poster (see Section 2.3), which helps users discover each other.
Refer to caption An exponential growth curve, approximately doubling every month, starting at 55k in May 2023 and exceeding 1.3M in October 2023.
Figure 2. Number of registered users on Bluesky since April 2023.
Bluesky launched an invite-only beta release in February 2023, and has grown to over 3 million registered users in January 2024, as shown in Figure 2. Bluesky Social PBC (a public-benefit corporation) develops the official client app and operates the core services; the client and several server-side components are open source under the MIT license (Bluesky Social PBC, [n. d.]c). The protocols they use are defined by open specifications (Bluesky Social PBC, [n. d.]b). Several parts of the system, such as feed generators (Section 2.3) and various alternative clients (Bluesky Social PBC, [n. d.]a) are developed and operated by independent third parties.
@@ 2.1. Moderation Features
Automated systems label potentially problematic content (such as images of a sexual or violent nature, posts promoting hate groups, or spam), and the app’s preferences allow users to choose whether to show or hide content in each of these categories in their feeds.
@@
One user can block another, which prevents all future interactions (such as mentions, replies, or reposts) between those accounts in addition to muting. Similarly to mutelists, a user can also publish a list of accounts, and other users can block all accounts on that list by subscribing to it.
@@
A user who makes a post can restrict who is allowed to reply to it (anyone, anyone they follow, anyone mentioned in the post, and/or anyone on a particular list of accounts) (Bluesky Social PBC, 2023c).
@@
Users can report content that violates the terms of service to server operators, and the operators can take down violating media, posts, or accounts.
@@ 2.2. User Handles
Like on Twitter/X, a Bluesky user has two names: the display name can be almost any string, and the handle needs to uniquely identify a user. A handle, prefixed with an @ sign, is used to mention another user in a post. Examples can be seen in Figure 1 (the display name is in bold, and the handle is in a smaller font and lighter color).
The need for handles to be unique creates challenges in decentralized systems, since it requires an authority that determines which handle is assigned to which user. Mastodon’s approach is to include the server name in the handle, which makes it difficult to move to another server. An alternative would be to use a blockchain-based naming system, such as the Ethereum Name System (ENS) (ENS Labs Limited, [n. d.]); this has the disadvantage of requiring the user to buy cryptocurrency in order to create an account, which we wanted to avoid.
Instead, Bluesky and atproto use DNS domain names as handles. If a user already owns a domain name, they can claim it as their Bluesky handle by adding a DNS record or by hosting a file under a /.well-known/ HTTPS URL on that domain (Liu, 2023). Users can also buy a new domain name within Bluesky, via a partnership with a domain registrar (Bluesky Social PBC, 2023g). Alternatively, users can sign up for a subdomain of .bsky.social for free.
Using DNS domain names as handles has several advantages:
@@
  • Users do not need to host their own server to use their own domain name; a DNS record requires only a one-time setup and no ongoing maintenance.
@@
  • An organization can easily allow their staff to demonstrate their affiliation by granting them handles that are subdomains of the organization’s main domain name (comparable to institutional email addresses). For example, a journalist’s handle may indicate that they are at a particular news organization.
  • Providers wanting to offer free subdomains can do so at very little cost.
@@ 2.3. Custom Feeds and Algorithmic Choice
Several decentralized social networks choose to offer only a reverse-chronological feed of posts from accounts the user is following – a backlash against the opaque content recommendation algorithms employed by mainstream centralized social networks. For example, Mastodon advertises itself as having “no algorithms or ads to waste your time” (Mastodon gGmbH, [n. d.]).
Our belief is that the problem lies not with algorithms per se, but rather with centrally controlled, opaque algorithms that remove user agency and prioritize user engagement over all else, e.g. by promoting controversial posts. Good recommendation algorithms can help users discover content that is relevant to them and find new accounts to follow – especially important for new users who are not yet following many accounts. They are also helpful for surfacing content on a particular topic, whereas following a user means seeing all of their posts, which might be on a mixture of topics, not all necessarily interesting to all followers. Giving users the ability to choose their algorithms lets them control what they want to see, rather than having the platform decide for them.
Bluesky Social PBC offers a selection of feed algorithms of its own, and also allows anybody to create their own feed generator (Bluesky Social PBC, 2023a). Tens of thousands of custom feeds have already been created. Our goal is to offer an open and diverse marketplace of algorithms in which communities can adapt the system to suit their needs, and users have more agency over how they spend their time and attention (Graber, 2023). Section 3.4 explains how feed generators work.
In Figure 1, a selection of bookmarked feeds is given at the top of the screen; in this example, the selected “Following” feed is the default reverse-chronological timeline, while “Week Peak Feed” (network-wide posts with many likes from the last week) and “Birds!” (photos and posts from birdwatchers) are third-party feeds. A feed generator can use arbitrary criteria to select its content. For example, the birdwatching feed uses a manually curated list of accounts, and selects posts from those accounts that contain a #birds hashtag or a feather emoji character. Alternative approaches, such as machine learning algorithms, are equally possible.
@@ 3. The AT Protocol Architecture
The data types and concepts for a particular social mode are defined by a lexicon, which specifies the schema of the data and the request endpoints involved in providing that social mode (Bluesky Social PBC, [n. d.]b). At the moment, the com.atproto lexicon defines the core AT Protocol concepts such as user identity (Section 3.5), and the app.bsky lexicon defines the microblogging mode. Anyone can define a new lexicon, allowing new social modes to coexist alongside the Bluesky social app on a shared infrastructure. The purpose of a lexicon is to provide documentation, to allow code generation and type-checking in applications, and to facilitate the process of specifying and versioning an interoperable protocol.
The biggest constraint for new social modes is that atproto is currently designed for content that users want to make publicly available. In particular, Bluesky user profiles, posts, follows, and likes are all public. Blocking actions are also currently public; however, we are investigating mechanisms for making these private (Bluesky Social PBC, 2023h; Newbold, 2023). At present, Bluesky does not support private communications, such as direct messages, though we plan to add this in the future. Only a small amount of user state is currently private: any muted accounts and threads, notifications and their read/unread status, and user preferences such as pinned feeds and content filtering settings.
@@ 3.1. User Data Repositories
All data that a user wishes to publish is added to their repository, which stores a collection of records. Whenever a user performs some action – making a post, liking another user’s post, following another user, etc. – that action becomes a record in their repository. Records are encoded in DAG-CBOR (Protocol Labs, [n. d.]), a restricted form of CBOR (Bormann and Hoffman, 2020), a compact binary data format. The schema of records is defined by the lexicon, and a repository may contain a mixture of records from several different lexicons, representing user actions in different social modes. Media files (e.g. images) are stored outside of the user’s repository, but referenced by their CID (IPFS Documentation, [n. d.]) (essentially a cryptographic hash) from a record in the repository. Similarly, a reference to a record in another repository (e.g. identifying a post being liked) also includes its CID.
@@ 3.2. Personal Data Servers (PDS)
Compared to choosing a Mastodon server, the user’s choice of PDS hosting provider is fairly inconsequential. The PDS URL is internal to the system, and is not normally visible to users. It makes no difference whether two users are on the same PDS or different PDSes, since interaction between users goes via the indexing infrastructure in any case. A user can migrate from one PDS to another by simply copying their repository and media files to the new PDS, and pointing their account ID at the new PDS URL (see Section 3.5). Even if a PDS shuts down without warning, users can upload a backup of their repository to a new PDS, and thus recover their account without losing any of their posts or their social graph.
PDS operators will generally want to perform some basic moderation by deleting any illegal content hosted on their servers. However, PDS-level moderation is much less important than server-level moderation in Mastodon, because in atproto, the primary moderation role is taken on by seperate actors in the system – the labelers and feed generators (see Section 3.4). This allows different sets of people to offer server hosting and moderation services, respectively; we believe this separation is valuable since operating a server and moderating a community require largely disjoint sets of skills (Roth and Lai, 2023).
At the time of writing, Bluesky’s indexing infrastructure (see Section 3.3) only indexes repositories on PDS instances hosted by Bluesky Social PBC itself; this limitation exists to limit infrastructure load and abuse problems during the beta period. In that sense, Bluesky is not yet fully decentralized. Support for third-party PDS operators is already implemented and enabled in Bluesky’s sandbox (testing) environment, and a PDS implementation suitable for self-hosting is already open source (Bluesky Social PBC, [n. d.]c). We plan for the Bluesky indexing infrastructure to begin indexing repositories on other PDS operators (indicated by dashed arrows in Figure 3) in early 2024.
@@ 3.3. Indexing Infrastructure
At the time of writing, most of Bluesky’s indexing infrastructure is operated by Bluesky Social PBC (indicated by a shaded area in Figure 3). However, the company does not have any privileged access: since repositories are public, anybody can crawl and index them using the same protocols as our systems use. Client apps can switch to reading from a different index, or use a combination of multiple indexes.
While operating a small PDS is designed to be cheap, operating an indexer that ingests the entire network requires greater computing resources. We therefore expect that there will be fewer hobbyist indexers than self-hosted PDSes. Nevertheless, as Bluesky grows, there are likely to be multiple professionally-run indexers for various purposes. For example, a company that performs sentiment analysis on social media activity about brands could easily create a whole-network index that provides insights to their clients. Web search engines can incorporate Bluesky activity into their indexes, and archivists such as the Internet Archive can preserve the activity for posterity.
@@ 3.3.1. The Relay
The first component is the Relay, which crawls the user repositories on all known PDSes and consumes the streams of updates that they produce. The Relay checks the signatures and Merkle tree proofs on updates, and maintains its own replica of each repository. From this information, the Relay creates the firehose: an aggregated stream of updates that notifies subscribers whenever records are added or deleted in any of the known repositories.
The firehose is publicly available. Consuming the firehose is an easier way of building an index over the whole network, compared to directly subscribing to the source PDSes, since the Relay performs some initial data cleaning such as discarding malformed updates and filtering out high-volume spam. The firehose can optionally include Merkle proofs and signatures along with records, allowing subscribers to check that they are authentic.

### v2

@@ Abstract.
Bluesky is a new social network built upon the AT Protocol, a decentralized foundation for public social media. It was launched in private beta in February 2023, and has grown to over 10 million registered users by October 2024. In this paper we introduce the architecture of Bluesky and the AT Protocol, and explain how the technical design of Bluesky is informed by our goals: to enable decentralization by having multiple interoperable providers for every part of the system; to make it easy for users to switch providers; to give users agency over the content they see; and to provide a simple user experience that does not burden users with complexity arising from the system’s decentralized nature. The system’s openness allows anybody to contribute to content moderation and community management, and we invite the research community to use Bluesky as a dataset and testing ground for new approaches in social media moderation.
social networks; decentralization; federation
journalyear: 2024copyright: ccconference: Proceedings of the ACM Conext-2024 Workshop on the Decentralization of the Internet; December 9–12, 2024; Los Angeles, CA, USAbooktitle: Proceedings of the ACM Conext-2024 Workshop on the Decentralization of the Internet (DIN ’24), December 9–12, 2024, Los Angeles, CA, USAdoi: 10.1145/3694809.3700740isbn: 979-8-4007-1252-4/24/12ccs: Information systems Social networking sitesccs: Social and professional topics Centralization / decentralizationccs: Networks Social media networks
@@ 1. Introduction
Over the last two decades, social media has evolved into a “digital town square” and a cornerstone of civic life (Barabas et al., 2017). When a social media service is under the full control of a single corporation, it may change its policies on the whim of its leaders (Yeung, 2023). Its operations are opaque (e.g. regarding which content is recommended to users), and its users lack agency over their user experience. As a result, there has been increasing interest in decentralized social networks, of which the fediverse around the ActivityPub protocol (Lemmer-Webber et al., 2018) and the Mastodon software (Mastodon gGmbH, 2024) is perhaps the best known.
@@
In this paper we introduce the AT Protocol (atproto), a decentralized foundation for social networking, and Bluesky, a Twitter-style microblogging app built upon it. A core design goal of atproto and Bluesky is to enable a user experience of the same or better quality as centralized services, while being open and decentralized on a technical level. The AT Protocol is designed such that for every part of the system there can be multiple competing operators providing interoperable services, making it easy to switch providers.
Decentralization alone is not able to solve some of the thorniest problems of social media, such as misinformation, harassment, and hate speech (Roth and Lai, 2023). However, by opening up the internals of a service to the public, decentralization can enable a marketplace of approaches to these problems (Masnick, 2019). For example, Bluesky allows anybody to run moderation services that make subjective decisions of selecting desirable content or flagging undesirable content, and users can choose which moderation services they want to subscribe to. Moderation services are decoupled from hosting providers, making it easy for users to switch moderation services until they find ones that match their preferences. We hope that this architectural openness enables communities to develop their own approaches to managing problematic content, independently of what hosting providers implement (Masnick, 2019).
For example, researchers wanting to identify disinformation campaigns can easily get access to all content being posted, the social graph, and user profiles on Bluesky. If they are able construct an algorithm to label suspected disinformation, they can publish their labels in real time, and users who wish to see those labels can enable them in their client. One goal of this paper is to bring Bluesky and the AT Protocol to the attention of researchers working on such algorithms, and to invite them to use the rapidly growing dataset of Bluesky content as a basis for their work.
@@ 2. The Bluesky Social App
Bluesky is a microblogging application in the style of Twitter/X. The “official” client app is available on iOS, Android, and the web; several independently developed client apps are also available, such as Graysky (Newman, 2024) and deck.blue (Filho, 2024). Users can make public posts containing up to 300 characters of text, up to four images, and short-form video. They can interact with posts by replying, reposting, or liking, and they can follow other users. By default users have two feeds: one showing posts by followed accounts in reverse chronological order, and one recommending popular content across the network. Users can also choose from many alternative feeds that show content on various topics, as explained below.
Bluesky launched an invite-only beta in February 2023, and has grown to over 10 million registered users in the following 20 months. Bluesky Social PBC (a public-benefit corporation) develops the official client app and operates the core services; the client and several server-side components are open source, dual-licensed under the MIT and Apache 2.0 licenses (Bluesky Social PBC, 2024d). The protocols they use are defined by open specifications (Bluesky Social PBC, 2024b). Several parts of the system, such as feed generators and alternative clients (Bluesky Social PBC, 2024a) are developed and operated by independent third parties.
@@ 2.1. Moderation Features
Automated systems label potentially problematic content (such as spam, or images of a sexual or violent nature), and the app’s preferences allow users to choose whether to show or hide content in each of these categories in their feeds. Some filters are always applied regardless of user preferences.
@@
One user can block another, which prevents all future interactions (such as mentions, replies, or reposts) between those accounts in addition to bidirectional muting. Similarly to mutelists, a user can also publish a list of accounts, and other users can block all accounts on that list by subscribing to it.
@@
A user who makes a post can restrict who is allowed to reply to it (anyone, anyone they follow, anyone mentioned in the post, and/or anyone on a particular list of accounts) (Bluesky Social PBC, 2023c). The original poster can also hide replies to be accessible but not shown by default.
Quote detachment::
When a post is quoted by another user’s post, the author of the original post can detach it from the quote post. This reduces the risk of harassment through dog-piling.
@@
Users can report content that violates the terms of service to moderation services. Operators of the services described in Section 3 can take down violating media, posts, or accounts.
@@ 2.2. User Handles
Like on Twitter/X, a Bluesky user has two names: the display name can be almost any string, and the handle needs to uniquely identify a user. A handle, prefixed with an @ sign, is used to mention another user in a post.
The need for handles to be unique creates challenges in decentralized systems, since it requires an authority that determines which handle is assigned to which user. Mastodon’s approach is to include the server name in the handle, which makes it difficult to move to another server. An alternative would be to use a blockchain-based naming system, such as the Ethereum Name System (ENS) (ENS Labs Limited, 2024); this has the disadvantage of requiring the user to buy cryptocurrency in order to create an account, which we wanted to avoid.
Instead, Bluesky and atproto use DNS domain names as handles. Users can sign up for a subdomain of .bsky.social for free. If a user already owns a domain name, they can claim it as their Bluesky handle by adding a DNS record or by hosting a file under a /.well-known/ HTTPS URL on that domain (Liu, 2023). Users can also buy a new domain name within Bluesky, via a partnership with a domain registrar (Bluesky Social PBC, 2023g). Domain names have several advantages:
@@
  • Users do not need to host their own server to use their own domain name; they only need to maintain a DNS record.
@@
  • An organization can allow its staff to demonstrate their affiliation by granting them subdomains of the organization’s main domain name (comparable to institutional email addresses). For example, a journalist’s handle may indicate that they are at a particular news organization.
  • Offering free subdomains costs very little.
@@ 2.3. Custom Feeds and Algorithmic Choice
Several decentralized social networks offer only a reverse-chronological feed of posts from accounts the user is following – a backlash against the opaque recommendation algorithms used by mainstream social networks. For example, Mastodon advertises itself as having “no algorithms or ads to waste your time” (Mastodon gGmbH, 2024).
Our belief is that the problem lies not with algorithms per se, but rather with centrally controlled, opaque algorithms that remove user agency and prioritize user engagement over all else, e.g. by promoting controversial posts. Good recommendation algorithms can help users discover content that is relevant to them and find new accounts to follow – especially important for new users who are not yet following many accounts. They are also helpful for surfacing content on a particular topic, whereas following a user means seeing all of their posts, which might be on a mixture of topics, not all necessarily interesting to all followers. Our goal is to offer an open and diverse marketplace of algorithms in which communities can adapt the system to suit their needs, and users have more agency over how they spend their time and attention (Graber, 2023).
Bluesky Social PBC offers a selection of feed algorithms of its own, and also allows anybody to create their own feed generator (Bluesky Social PBC, 2023a). Tens of thousands of custom feeds have already been created. A feed generator can use arbitrary criteria to select its content. For example, it may use a manually curated list of accounts, and select posts from those accounts that contain a particular hashtag or emoji character. Use of machine learning algorithms is equally possible.
@@ 3. The AT Protocol Architecture
A social mode is defined by a lexicon, which specifies the schema of the data and the request endpoints involved (Bluesky Social PBC, 2024b). Several lexicons are currently used: for example, the com.atproto lexicon defines the core AT Protocol concepts such as user identity (Section 3.5), the app.bsky lexicon defines the microblogging mode, and the tools.ozone lexicon supports moderation actions (Bluesky Social PBC, 2024c). Anyone can define a new lexicon, allowing new social modes to coexist alongside the Bluesky social app on a shared infrastructure. The purpose of a lexicon is to provide documentation, to allow code generation and type-checking in applications, and to facilitate schema evolution for long-term compatibility and interoperability.
The biggest constraint for new social modes is that atproto is currently designed for content that users want to make public. In particular, Bluesky user profiles, posts, follows, and likes are all public. Blocking actions are also currently public; we are investigating mechanisms for making these private (Bluesky Social PBC, 2023h; Newbold, 2023). Only a small amount of user state is currently private: muted accounts and threads, notifications and their unread status, and user preferences such as content filtering settings. Private communications (direct messages) are currently handled by a centralized service run by Bluesky Social PBC; we plan to decentralize this component in the future.
@@ 3.1. User Data Repositories
All data that a user wishes to publish is added to their repository, which stores a collection of records. Whenever a user performs some action – making a post, liking another user’s post, following another user, etc. – that action becomes a record in their repository. Records are encoded in DAG-CBOR (Protocol Labs, 2021), a restricted form of CBOR (Bormann and Hoffman, 2020), a compact binary data format. The schema of records is defined by the lexicon, and a repository may contain a mixture of records from several different lexicons, representing user actions in different social modes. Media files (e.g. images) are stored outside of the user’s repository, but referenced by their CID (IPFS Documentation, 2022) (essentially a cryptographic hash) from a record in the repository. Similarly, a reference to a record in another repository (e.g. identifying a post being liked) also includes its CID.
@@ 3.2. Personal Data Servers (PDS)
Compared to choosing a Mastodon server, the user’s choice of PDS hosting provider is fairly inconsequential. The PDS URL is internal to the system, and is not normally visible to users. It makes no difference whether two users are on the same PDS or different PDSes, since interaction between users goes via the indexing infrastructure in any case. A user can migrate from one PDS to another by copying their repository and media files to the new PDS, and pointing their account at the new PDS URL (Section 3.5). Even if a PDS shuts down without warning, users can upload a backup of their repository to a new PDS, and thus recover their account without losing any of their posts or their social graph.
PDS operators generally perform some basic moderation by deleting any illegal content hosted on their servers. However, PDS-level moderation is much less important than server-level moderation in Mastodon, because in atproto, the primary moderation role is taken on by separate actors in the system – the labelers and feed generators (see Section 3.4). This allows different sets of people to offer server hosting and moderation services, respectively; we believe this separation is valuable since operating a server and moderating a community require largely disjoint sets of skills (Roth and Lai, 2023).
@@ 3.3. Indexing Infrastructure
At the time of writing, most of Bluesky’s indexing infrastructure is operated by Bluesky Social PBC (indicated by a shaded area in Figure 1). However, the company does not have any privileged access: since repositories are public, anybody can crawl and index them using the same protocols as our systems use. If Bluesky Social PBC were to violate users’ expectations – for example, by censoring (omitting from the index) accounts that users wish to see – other parties would be free to provide their own indexes that do not perform this censorship. Client apps could switch to reading from a different index, or even use a combination of multiple indexes.
In July 2024 (when Bluesky had 6 million users), it was possible to maintain a real-time copy of all user repositories on a single server for US\$153 per month (Newbold, 2024a) – expensive for a hobby project, but easily affordable by many organizations. This figure includes repository storage and the inbound bandwidth for fetching records from PDSes, but not the computational resources to build and serve summary indexes from that data.
Due to the higher cost, we expect that there will be fewer hobbyist indexers than self-hosted PDSes. Nevertheless, as Bluesky grows, there are likely to be multiple professionally-run indexers for various purposes. For example, a company that performs sentiment analysis on social media activity about brands could easily create a whole-network index that provides insights to their clients. Web search engines can incorporate Bluesky activity into their indexes, and archivists such as the Internet Archive can preserve the activity for posterity.
@@ 3.3.1. The Relay
The first component is the Relay, which crawls the user repositories on all known PDSes and consumes the streams of updates that they produce. The Relay checks the signatures and Merkle tree proofs on updates, and maintains its own replica of each repository. From this information, the Relay creates the firehose: an aggregated stream of updates that notifies subscribers whenever records are added or deleted in any of the known repositories. Normally the WebSocket connection to PDSes provides the relay with low-latency notifications of repository changes, but in case of network interruptions the relay can also periodically re-crawl repositories and compare them to its local replica to determine what records have been added or deleted.
The firehose is publicly available. Consuming the firehose is an easier way of building an index over the whole network, compared to directly subscribing to the source PDSes, since the Relay performs some initial data cleaning (discarding malformed updates, filtering out illegal content and high-volume spam). The firehose includes Merkle proofs and signatures along with records, allowing subscribers to check that they are authentic.